# A Survey on Large Language Model based Human-Agent Systems

### v1

@@ 3.3.3 Coopetition
Coopetition is an interaction where cooperation and competition coexist at the same time. Within this interaction, participants collaborate on shared tasks or mutual goals while also seeking to outdo each other to improve their own performance or gain extra advantages. In terms of the LLM-HAS, this dual aspect implies that agents and human may join forces to address complex issues while competing in specific domains such as efficiency or precision. This approach not only combines the strengths of both collaboration and competition but also fosters innovation driven by competitive incentives while also reaping the benefits of cooperative synergy. Successfully managing coopetition typically requires mechanisms for building trust and adaptable strategies that reconcile collective advantages with personal aspirations, which is a challenge for the LLM-HAS. For example, humans and agents play the prisoner’s dilemm in the shared workspace (Jiang et al., 2025).
@@ 3.4 Orchestration Paradigm
The orchestration paradigm in LLM-HAS refers to how tasks and interactions are managed between humans and agents, covering two dimensions in our survey: Task Strategy (ordering) and Temporal Synchronization (timing).
@@ 3.4.1 Task Strategy
One-by-One. The one-by-one strategy requires participants (humans and LLM-based agents) to perform tasks sequentially, taking clearly defined turns. For example, a human first outlines a plan, the agent then executes the task, the human subsequently reviews the output, and finally, the agent refines its work based on feedback Chang et al. (2024); Zhou et al. (2025). Such sequential interaction helps maintain a clear order of execution and reduces the complexity associated with concurrent task management. However, this rigidity may limit overall efficiency and flexibility, especially in dynamic scenarios requiring parallel processing or rapid interaction cycles Bansal et al. (2024); Guo et al. (2024b).
Simultaneous. Simultaneous strategy describes an interaction pattern in which LLM-based agents and humans respond concurrently in real time. Compared to the one-by-one strategy, the simultaneous approach more closely mirrors real-world conditions encountered in many simulation tasks Liu et al. (2023b); Zhang et al. (2025). However, this strategy demands sophisticated mechanisms to handle latency mitigation and seamless coordination between participants.
3.4.2 Temporal Synchronization
@@
Synchronous (1) Real-time interaction: Humans and agents communicate simultaneously or in immediate sequence ; (2) Immediate response: Participants expect or require prompt feedback. (e.g. live chat session, real-time voice assistant).
@@ 3.5.1 Communication Structure
In Centralized structure, one primary agent or a group of core agents acts as a central node to coordinate all communications within the system. This central agent manages interactions among other agents, simplifying coordination and minimizing conflicts Cui et al. (2024). Decentralized structure employs peer-to-peer communication, enabling direct interactions among agents without centralized control. Agents autonomously manage their communications based on systemic information, enhancing system flexibility, adaptability, and robustness Shao et al. (2024); Xu et al. (2025). In addition, Hierarchical structure organizes agents into clearly defined levels, assigning distinct roles and responsibilities according to their position within the hierarchy Liu et al. (2023b); Pan et al. (2024c). High-level agents typically fulfill managerial or strategic roles, providing overarching guidance and supervision, while lower-level agents perform specialized tasks and execute detailed operations.
@@ 3.5.2 Communication Mode
Conversation. The conversation-based mode is perhaps the most prevalent and intuitive approach in LLM-HAS, wherein agents and humans directly engage through natural language dialogues. This interaction format typically utilizes conversational interfaces that allow iterative exchanges, questions, clarifications, and dynamic responses, facilitating efficient collaboration and mutual understanding Shao et al. (2024). For instance, conversational LLM agents can assist users by answering queries, explaining complex concepts, or collaboratively solving reasoning tasks through iterative dialogues Wang et al. (2024c). While intuitive and flexible, conversational interactions rely significantly on the communicative clarity and dialogue management capabilities of the LLM agents.
Observation. In the observation-based communication mode, agents acquire information implicitly by observing participants behaviors, decisions, or interactions within their environment, rather than through explicit verbal communication. This mode leverages indirect signals, including user actions, feedback cues, or behavioral traces, to infer intentions, preferences, or states Seo et al. (2025). For example, an LLM-driven tutoring system may adaptively provide targeted instructions by continuously observing student problem-solving behaviors without explicit verbal queries Pan et al. (2024c). However, relying solely on observational signals can introduce ambiguity, potentially impacting inference accuracy unless complemented by robust inferential mechanisms.
@@
MINT MINT (Wang et al., 2024c) Link
@@
MINT MINT (Wang et al., 2024c) Link
@@
Software Development MINT MINT (Wang et al., 2024c) Link
@@
4 Application
@@
Embodied AI. Embodied AI applications involve various aspects of dynamic and complex real-world tasks, benefiting from valuable humans’ feedback and interactions in Human-Agent collaboration for adaptation. Ye et al. (2023) explores incorporating LLMs in human-robotic collaboration assembly tasks, allowing seamless communication between robots and humans and increasing trust in human operators. To address the challenges of false planning due to suboptimal environment changes , Seo et al. (2025) proposes REVECA to enable efficient memory management and optimal planning. Additionally, Tanneberg et al. (2024b) extends the agents’ collaboration with a group of humans via Attentive Support, enabling agents’ ability to remain silent to not disturb the group if desired.
@@
Software Development. Given the inherently collaborative nature of software development, human-agent collaboration has emerged as a critical component in addressing the associated challenges. Feng et al. (2024) introduces ReHAC framework, wherein agents are trained to determine the optimal stages for human intervention within the problem solving process, offering improved generalizability over the traditional heuristic-based approaches. Building on this direction, Zhou et al. (2025); Han et al. (2025); Wang et al. (2024c) investigate broader spectrum of human feedback types via multi-turn human-agent interactions. These approaches incorporate carefully designed optimization objectives to effectively capture more diverse and nuanced interactions between humans and agents.
@@
Conversational Systems. Due to the frequent presence of ambiguity and the broad range of complex tasks in conversational systems, effective human-agent collaboration constitutes as a critical component of the system. Zhang et al. (2024c) introduces Proactive Agent Planning, wherein agents are trained to predict classification needs based on the user-agent conversational interactions and current environment, thereby leading to improved reasoning efficacy. Wu et al. (2022) introduces Chaining LLM to improve the quality of task outcomes and enhance the transparency, controllability and collaboration from the conversational systems.
@@
Gaming. Human-Agent collaboration is naturally well-suited to simulated gaming environments due to their dynamicity and sophistication. Such collaborative interactions have been shown to enhance humans’ experience, satisfaction and comprehension of both the environment and agents Gong et al. (2023); Gao et al. (2024c). Concurrently, these interactions also contribute to improved agents’ task performance and decision-making capabilities. For instance, MindAgent framework Gong et al. (2023) illustrates the efficacy of human-agent collaboration through measurable improvements in task outcomes when humans and agents work together. Mehta et al. (2024) demonstrates agents achieve improved outcomes when interacting with humans via autonomous confusion detection and clarification questions inquiries. Ait et al. (2024) introduces Meta-Command Communication-based framework to enable effective human-agent collaboration. To address challenges related to execution latency while maintaining strong reasoning capabilities, Liu et al. (2023a) proposes Hierarchical Language Agent that promotes faster responses, stronger cooperation, and more consistent language communications.
@@
Finance. Given the inherent complexity of stock markets and financial data systems, where investors’ strategies and risk preferences are critical determinants of successful outcomes, Human-Agent collaboration is increasingly recognized as an essential component in financial decision-making. FinArena Xu et al. (2025) has been proven effective in stock predictions by integrating the dynamic yet essential collaboration between experienced investors and advanced AI Agents. This collaborative framework is demonstrated to produce optimal investment outcomes in terms of the best annualized return and the sharpe ratio for investors Xu et al. (2025).
@@ 5.1 Human-Agent Framework
In this section, we present a comprehensive introduction to the three open-source LLM-HAS frameworks from previous works: Collaborative Gym Shao et al. (2024), COWPILOT Huq et al. (2025), and DPT-Agent Zhang et al. (2025). Although all three employ an LLM-HAS architecture, they differ in key configuration aspects, including environment settings, interaction types, orchestration paradigms, and communication strategies. Collaborative Gym Shao et al. (2024) facilitates asynchronous interactions among humans, agents, and task environments, supporting various simulated and real-world tasks such as travel planning, data analysis, and academic writing. It emphasizes flexible, real-time collaboration and evaluates both outcomes and interaction quality, making it a robust tool for studying human-agent dynamics. COWPILOT Huq et al. (2025) provides a framework for human-agent collaborative web navigation through a Chrome extension. It employs a "Suggest-then-Execute" model under human supervision, allowing dynamic interventions to enhance task completion rates and reduce human workload. COWPILOT effectively demonstrates how human intervention can significantly improve agent performance. DPT-Agent Zhang et al. (2025) applies Dual Process Theory (DPT) to enable real-time simultaneous human-agent interactions. It features intuitive, fast decision-making and deliberative reasoning components, employing Theory of Mind and asynchronous reflection to manage latency and adapt dynamically to human actions. This approach excels in environments requiring immediate and adaptive responses.
Other notable frameworks, such as A2C Tariq et al. (2025), FinArena Xu et al. (2025), and a human-robot collaboration framework Liu et al. (2023a), also contribute significantly to specific domains like cybersecurity, financial forecasting, and robotic manipulation, respectively. These frameworks further demonstrate the diverse potential and adaptability of LLM-HAS.
@@ 5.2 Datasets and Benchmarks
We summarize the commonly used datasets and benchmarks for Large Language Model-based Human-Agent Systems in Table 3. Diverse domains employ distinct methodologies for evaluating these systems, aligned closely with their unique application contexts. Within the domain of embodied AI, the primary approach involves simulated environments Sun et al. (2024b, a); Mehta et al. (2024), designed to assess how effectively agents cooperate and execute tasks in dynamic, interactive scenarios. Another significant domain, Conversational Systems, encompasses applications such as question answering Feng et al. (2024), website navigation Lù et al. (2024), design decision assistance Sharma et al. (2024), and travel planning Zhang et al. (2024c), adopting benchmarks that evaluate the ability of language models to function as user-aligned conversational assistants, ensuring interactions meet user expectations. Despite the extensive application coverage of current benchmarks, there remains a clear necessity for the development of more comprehensive and standardized benchmarking frameworks.
@@ 6 Challenges and Opportunities
LLM-HAS is designed to improve solutions for daily tasks and complex challenges like advanced reasoning. By integrating the intelligence of humans and LLM-based agents, tasks can be solved wisely and efficiently. However, implementing LLM-HAS may also bring out the duality of transformative potential and significant risk. On one hand, the remarkable capabilities of LLM agent in natural language understanding, generation Zou et al. (2024a, b), and emergent reasoning Wang et al. (2024b); Gu et al. (2025) have catalyzed their integration into increasingly sophisticated agentic systems Xi et al. (2025). However, this potential is counterbalanced by several fundamental challenges. These challenges can be divided into five distinct aspects: (1) Mostly Agent-Centered Work (2) Human Flexibility and Variability (3) Inadequate Evaluation Methodologies (4) Unresolved Safety Vulnerabilities (5) Fine-Grained Collaboration Type.
Mostly Agent-Centered Work. In most LLM-HAS studies, guidance flows in usually in a single direction, with humans evaluating agent outputs and providing corrective or evaluative feedback. Namely, the current studies are mostly agent-centered. However, enabling agents to observe human actions, detect errors or inefficiencies, and offer timely suggestions can transform collaboration and reduce human effort by leveraging agent intelligence. When agents act as instructors by proposing alternative strategies, drawing attention to overlooked risks, and reinforcing effective practices as tasks unfold in real time, both the human and agent benefit. Genuine collaboration arises when humans and LLM-based agents stand as equal partners and give equal weight to each other’s insights. However, current work primarily focuses on delegation rather than coordination or cooperation, leaving significant potential for feedback loops driven by agents. We believe that shifting toward human-centered system, or an equalized LLM-HAS, will unlock the full promise of teamwork between humans and agents.
@@
Human Flexibility and Variability. Human feedback varies widely in terms of role, timing, and style across various LLM-HAS. Human are usually subjective based on their personalities. Namely, different humans in LLM-HAS may lead to different outcomes and conclusions. In addition, humans, regarded as a "special agent" in the LLM-HAS, are subject to fewer restrictions and evaluations than LLM-based agents. This limits how the LLM-HAS can be improved because the impedance may be on the human side instead of the agent. This concern remains and requires a refined strategy to define the strict, fine interaction rule and evaluation equally for both human and LLM-based agents. Also, many studies today substitute real human participants with LLM simulated human proxies, failing to capture human input’s variety and unpredictability. The performance gap between humans and the simulated human remains unknown, potentially making the comparison incomparable.
@@
Inadequate Evaluation Methodologies. In existing evaluation frameworks for LLM-HAS, improvements focus primarily on agent accuracy and static benchmarks, which ignore the real burden placed on human collaborators. People dedicate varying amounts of time, attention and cognitive effort depending on the type and frequency of feedback they must provide, yet no standard metric captures this human workload or its impact on overall efficiency. Evaluation methods should measure factors such as time spent offering feedback, perceived mental workload and effort required to detect and correct errors, and they should cover every phase of the human agent collaboration from initial task assignment through post execution review. As human expertise and LLM based agent capabilities merge to deliver unprecedented performance, both uncertainty and variability grow. A new evaluation approach or set of metrics that systematically and comprehensively quantifies contributions and costs for both humans and agents is essential to ensure truly efficient collaboration.
@@
Unresolved Safety Vulnerabilities. In research on LLM-HAS the emphasis on improving agent performance has left safety, robustness and privacy underexplored in the context of human interaction. As people and LLM-based agents collaborate in dynamic workflows, the risk of misaligned behavior, unexpected failures, or unintended disclosure of sensitive information grows. Humans engaging with these systems need clear safeguards around data sharing, error recovery protocols when agents behave unpredictably and privacy protections that cover every stage of the interaction. Robustness measures must ensure agents handle ambiguous or adversarial inputs without passing harm on to their human partners. Without studies that foreground human experience in safety and privacy design, real-world deployments will struggle to gain trust or meet acceptable risk thresholds. Rigorous investigation of how safety, robustness and privacy shape human agent workflows from design through deployment is essential to build collaborations that are both effective and respectful of human needs.
Fine-Grained Collaboration Type. In current research on LLM-HAS, broad interaction categories such as collaboration or competition are typically clearly defined, but granular-level subtypes such as delegation, coordination, and supervision remain underspecified. Yet these subtypes play a critical role in shaping the division of labor, communication patterns, and decision-making protocols between humans and agents. For instance, delegation implies that a human issues a goal and the agent independently plans and executes the necessary steps. In contrast, coordination involves dynamic sharing of responsibilities and mutual adjustment of actions. Without a precise taxonomy of interaction subtypes, it is impossible to compare different systems, establish standard evaluation benchmarks or ensure predictable behavior. This ambiguity is particularly harmful when executing real-time or safety-critical tasks, where unclear role definitions may lead to miscommunication, operational failures, or reduced trust. Developing a systematic framework that clearly defines each interaction subtype will therefore enable rigorous experimental design, more reliable performance and smoother collaboration in complex LLM-HAS.

### v2

@@ 3.3.3 Coopetition
Coopetition is an interaction where cooperation and competition coexist at the same time. Within this interaction, participants collaborate on shared tasks or mutual goals while also seeking to outdo each other to improve their own performance or gain extra advantages. In terms of the LLM-HAS, this dual aspect implies that agents and human may join forces to address complex issues while competing in specific domains such as efficiency or precision. This approach not only combines the strengths of both collaboration and competition, but also fosters innovation driven by competitive incentives while also reaping the benefits of cooperative synergy. Successfully managing coopetition typically requires mechanisms for building trust and adaptable strategies that reconcile collective advantages with personal aspirations, which is a challenge for the LLM-HAS. For example, humans and agents play the prisoner’s dilemma in the shared workspace (Jiang et al., 2025).
@@ 3.4 Orchestration Paradigm
The orchestration paradigm in LLM-HAS refers to how tasks and interactions are managed between humans and agents, covering two dimensions in our survey: Task Strategy (ordering) and Temporal Synchronization (timing). Table 2 summarizes the two dimensions of the orchestration paradigm.
@@ 3.4.1 Task Strategy
Synchronous (1) Real-time interaction: Humans and agents communicate simultaneously or in immediate sequence; (2) Immediate response: Participants expect or require prompt feedback. (e.g., live chat session, real-time voice assistant).
@@
One-by-One. The one-by-one strategy requires participants (humans and LLM-based agents) to perform tasks sequentially, taking clearly defined turns. For example, a human first outlines a plan, the agent then executes the task, the human subsequently reviews the output, and finally, the agent refines its work based on feedback Liu et al. (2024a); Zhou et al. (2025). Such sequential interaction helps maintain a clear order of execution and reduces the complexity associated with concurrent task management. However, this rigidity may limit overall efficiency and flexibility, especially in dynamic scenarios requiring parallel processing or rapid interaction cycles Bansal et al. (2024); Guo et al. (2024b).
Simultaneous. Simultaneous strategy describes an interaction pattern in which LLM-based agents and humans respond concurrently in real time. Compared to the one-by-one strategy, the simultaneous approach more closely mirrors real-world conditions encountered in many simulation tasks Wang et al. (2025d); Zhang et al. (2025). However, this strategy demands sophisticated mechanisms to handle latency mitigation and seamless coordination between participants.
3.4.2 Temporal Synchronization
@@ 3.5.1 Communication Structure
In Centralized structure, one primary agent or a group of core agents acts as a central node to coordinate all communications within the system. This central agent manages interactions among other agents, simplifying coordination and minimizing conflicts Cui et al. (2024). Decentralized structure employs peer-to-peer communication, enabling direct interactions among agents without centralized control. Agents autonomously manage their communications based on systemic information, enhancing system flexibility, adaptability, and robustness Shao et al. (2024); Driess et al. (2023). In addition, Hierarchical structure organizes agents into clearly defined levels, assigning distinct roles and responsibilities according to their position within the hierarchy Liu et al. (2023b); Pan et al. (2024b). High-level agents typically fulfill managerial or strategic roles, providing overarching guidance and supervision, while lower-level agents perform specialized tasks and execute detailed operations.
@@ 3.5.2 Communication Mode
Conversation. The conversation-based mode is currently the most prevalent and intuitive approach in LLM-HAS, wherein agents and humans directly engage through natural language dialogues. This interaction format typically utilizes conversational interfaces that allow iterative exchanges, questions, clarifications, and dynamic responses, facilitating efficient collaboration and mutual understanding Shao et al. (2024). For instance, conversational LLM agents can assist users by answering queries, explaining complex concepts, or collaboratively solving reasoning tasks through iterative dialogues Wang et al. (2024b). While intuitive and flexible, conversational interactions rely significantly on the communicative clarity and dialogue management capabilities of LLM agents.
Observation. In the observation-based communication mode, agents acquire information implicitly by observing participants behaviors, decisions, or interactions within their environment, rather than through explicit verbal communication. This mode leverages indirect signals, including user actions, feedback cues, or behavioral traces, to infer intentions, preferences, or states Lu et al. (2024). For example, an LLM-driven tutoring system may adaptively provide targeted instructions by continuously observing student problem-solving behaviors without explicit verbal queries Pan et al. (2024b). However, relying solely on observational signals can introduce ambiguity, potentially impacting inference accuracy unless complemented by robust inferential mechanisms.
@@
MINT MINT (Wang et al., 2024b) Link
@@
MINT MINT (Wang et al., 2024b) Link
@@
Software Development MINT MINT (Wang et al., 2024b) Link
@@
4 Applications
A diverse range of applications has emerged for LLM-HAS. We elaborate on the five most frequent domains below and summarize corresponding datasets and benchmarks in Table 3. With new applications appearing almost weekly in this fast-growing field, we maintain an open-source GitHub repository to track recent developments.
@@ 4 Applications
Embodied AI. Applications in Embodied AI involve various aspects of dynamic and complex real-world tasks, benefiting from valuable human feedback and interactions in LLM-HAS. Ye et al. (2023) explores incorporating LLMs in human-robotic collaboration assembly tasks, allowing seamless communication between robots and humans and increasing trust in human operators. To address the challenges of false planning due to suboptimal environment changes, Seo et al. (2025) proposes REVECA to enable efficient memory management and optimal planning. Additionally, Tanneberg et al. (2024b) extends the agents’ collaboration with a group of humans via Attentive Support, enabling agents’ ability to remain silent to not disturb the group if desired.
@@
Software Development. The inherently collaborative nature of software development makes human-agent collaboration vital to improve development efficiency Lu et al. (2024); Han et al. (2025); Zhou et al. (2025). Feng et al. (2024) introduces ReHAC framework, wherein agents are trained to determine the optimal stages for human intervention within the problem-solving process, offering improved generalizability over the traditional heuristic-based approaches. Building on this direction, Zhou et al. (2025); Han et al. (2025); Wang et al. (2024b) investigate a broader spectrum of human feedback types via multi-turn human-agent interactions. These approaches incorporate carefully designed optimization objectives to effectively capture more diverse and nuanced interactions between humans and agents.
@@
Conversational Systems. In conversational systems, due to the frequent presence of ambiguity and lack of necessary information that agents cannot reliably infer, such as login credentials and payment details, effective human-agent collaboration constitutes a critical component of the system. Zhang et al. (2024c) introduces Proactive Agent Planning, wherein agents are trained to predict classification needs based on the user-agent conversational interactions and current environment, thereby leading to improved reasoning efficacy. Wu et al. (2022) introduces Chaining the LLM to improve the quality of task outcomes and enhance the transparency and controllability of the conversational systems.
@@
Gaming. LLM-HAS are naturally well-suited to simulated gaming environments due to their dynamicity and sophistication. Proper human-agent interactions have been shown to enhance humans’ experience, satisfaction and understanding of both the environment and agents Gong et al. (2023); Gao et al. (2024c). Collaborative interactions also contribute to improved agents’ task performance and decision-making capabilities. For instance, MindAgent framework Gong et al. (2023) illustrates the efficacy of human-agent collaboration through measurable improvements in task outcomes when humans and agents work together. Mehta et al. (2024) demonstrates agents achieve improved outcomes when interacting with humans via autonomous confusion detection and clarification questions and inquiries. Ait et al. (2024) introduces Meta-Command Communication-based framework to enable effective human-agent collaboration. To address challenges related to execution latency while maintaining strong reasoning capabilities, Liu et al. (2023a) proposes Hierarchical Language Agent that promotes faster responses, stronger cooperation, and more consistent language communications.
@@
Finance. Given the complexity of stock markets and financial systems, where investors’ strategies and risk preferences are critical determinants of outcomes, human-agent collaboration is increasingly recognized as a valuable paradigm. FinArena Xu et al. (2025) demonstrates the potential of integrating experienced investors with advanced AI agents to support stock prediction tasks. This collaborative framework has been shown to improve investment performance, yielding competitive annualized returns and Sharpe ratios Xu et al. (2025).
@@ 5.1 Human-Agent Framework
This section provides a detailed introduction to three representative open-source LLM-HAS frameworks: Collaborative Gym Shao et al. (2024), COWPILOT Huq et al. (2025), and DPT-Agent Zhang et al. (2025). They differ in key configuration aspects, including environment settings, interaction types, orchestration paradigms, and communication strategies. Specifically, Collaborative Gym Shao et al. (2024) facilitates asynchronous interactions among humans, agents, and task environments, supporting various simulated and real-world tasks such as travel planning, data analysis, and academic writing. It emphasizes flexible, real-time collaboration and evaluates both outcomes and interaction quality, making it a robust tool for studying human-agent dynamics. COWPILOT Huq et al. (2025) provides a framework for human-agent collaborative web navigation through a Chrome extension. It employs a "Suggest-then-Execute" model under human supervision, allowing dynamic interventions to enhance task completion rates and reduce human workload. It effectively demonstrates how human intervention can significantly improve agent performance. DPT-Agent Zhang et al. (2025) applies Dual Process Theory (DPT) to enable real-time simultaneous human-agent interactions. It features intuitive, fast decision-making and deliberative reasoning components, employing Theory of Mind and asynchronous reflection to manage latency and adapt dynamically to human actions. This approach excels in environments requiring immediate and adaptive responses.
Other frameworks, such as A2C Tariq et al. (2025), FinArena Xu et al. (2025), and human-robot collaboration framework Liu et al. (2023a), also contribute significantly to specific domains like cybersecurity, financial forecasting, and robotic manipulation, respectively. These frameworks further demonstrate the diverse potential and adaptability of LLM-HAS.
@@ 5.2 Datasets and Benchmarks
We summarize the commonly used datasets and benchmarks for Large Language Model-based Human-Agent Systems in Table 3. Diverse domains employ distinct methodologies for evaluating these systems, aligned closely with their unique application contexts. Within the domain of embodied AI, the primary approach involves simulated environments Sun et al. (2024b, a); Mehta et al. (2024), designed to assess how effectively agents cooperate and execute tasks in dynamic, interactive scenarios. Another significant domain, Conversational Systems, encompasses applications such as question answering Feng et al. (2024), website navigation Lù et al. (2024); Levy et al. (2024), design decision assistance Sharma et al. (2024), and travel planning Zhang et al. (2024c), adopting benchmarks that evaluate the ability of language models to function as user-aligned conversational assistants, ensuring interactions meet user expectations. Despite the extensive application coverage of current benchmarks, there remains a clear necessity for the development of more comprehensive and standardized benchmarking frameworks.
@@ 6 Challenges and Opportunities
In this section, we highlight some existing challenges and opportunities for LLM-HAS.
@@
Human Flexibility and Variability. Human feedback varies widely in terms of role, timing, and style across various LLM-HAS. Humans are often subjective, influenced by their personalities, which means different individuals interacting with an LLM-HAS may lead to different outcomes and conclusions. This highlights the need and opportunity for i) thorough investigations or benchmarks on how varied human feedback affects entire systems, and ii) flexible frameworks that can support and adapt to diverse human feedback. In addition, humans, regarded as a “special agent” in the LLM-HAS, are subject to fewer restrictions and evaluations than LLM-based agents. This limits how the LLM-HAS can be improved because the impedance may be on the human side instead of the agent. This concern remains and requires a refined strategy to define the strict, fine interaction rule and evaluation equally for both human and LLM-based agents. Also, many studies today substitute real human participants with LLM simulated human proxies, failing to capture human input’s variety and unpredictability. The performance gap between humans and the simulated human remains unknown, potentially making the comparison incomparable.
Mostly Agent-Centered Work. In most LLM-HAS studies, guidance flows in a single direction, with humans evaluating agent outputs and providing corrective or evaluative feedback. Namely, the current studies are mostly agent-centered. However, enabling agents to observe human actions, detect errors or inefficiencies, and offer timely suggestions can transform collaboration and reduce human effort by leveraging agent intelligence. When agents act as instructors by proposing alternative strategies, drawing attention to overlooked risks, and reinforcing effective practices as tasks unfold in real time, both humans and agents benefit. We believe that exploring human-centered LLM-HAS, or shifting toward an equalized LLM-HAS, will unlock the full promise of teamwork between humans and agents.
@@
Inadequate Evaluation Methodologies. In existing evaluation frameworks for LLM-HAS, improvements focus primarily on agent accuracy and static benchmarks, which ignore the real burden placed on human collaborators Ma et al. (2025). People dedicate varying amounts of time, attention and cognitive effort depending on the type and frequency of feedback they must provide, yet no standard metric captures this human workload or its impact on overall efficiency. Evaluation methods should measure factors such as time spent offering feedback, perceived mental workload and effort required to detect and correct errors, and they should cover every phase of the human agent collaboration from initial task assignment through post execution review. As human expertise and LLM-based agent capabilities merge to deliver unprecedented performance, both uncertainty and variability grow. A new evaluation approach or set of metrics that systematically and comprehensively quantifies contributions and costs for both humans and agents is essential to ensure truly efficient collaboration.
@@
Unresolved Safety Vulnerabilities. Most LLM-HAS works emphasize improving agent performance and have left safety, robustness and privacy underexplored in the context of human interaction Qiu et al. (2025). As people and LLM-based agents collaborate in dynamic workflows, the risk of misaligned behavior, unexpected failures, or unintended disclosure of sensitive information grows. Humans engaging with these systems need clear safeguards around data sharing, error recovery protocols when agents behave unpredictably and privacy protections that cover every stage of the interaction. Robustness measures must ensure agents handle ambiguous or adversarial inputs without passing harm on to their human partners Glickman and Sharot (2025). Without studies that emphasize human experience in safety and privacy design, real-world deployments will struggle to gain trust or meet acceptable risk thresholds. Rigorous investigation of how safety, robustness and privacy shape human agent workflows from design through deployment is essential to build collaborations that are both effective and respectful of human needs.
Applications and Beyond. The potential of LLM-HAS extends well beyond current applications. Many opportunities remain to be explored in challenging domains such as healthcare, finance, scientific research, education, and so on. While fully autonomous LLM-based agent systems encounter difficulties in handling complex, long-term tasks and earning full trust in safety and reliability, the involvement of humans to provide additional information, feedback, and control allows LLM-HAS to greatly improve overall system performance and safety. This opens the door to impactful applications across a broad range of critical fields.